# WonderHuman: Hallucinating Unseen Parts in Dynamic 3D Human Reconstruction

### v1

@@ 5.3.1 ZJU-Mocap(revised), MVHumanNet, and Monocap Datasets
These three datasets serve as the primary testbeds for our experiments due to the availability of ground truths for invisible parts. The quantitative results are presented in Tab. I, where our method surpasses all evaluation techniques in all three metrics, indicating its efficacy in reconstructing both geometry and appearance for invisible parts. Qualitatively, as shown in Fig. 4, the limitations of all compared methods become more evident when visualizing invisible parts. Methods such as GaussianAvatar, ExAvatar, and SplattingAvatar, which are based on Gaussian Splatting, exhibit noticeable artifacts and inconsistencies, including noisy textures and blank spots. Instant-NVR and HumanNeRF, due to their NeRF-based ray-shooting geometric reconstruction technique, not only struggle with appearance consistency but also suffer from geometric issues like floating artifacts and penetrating holes, diminishing the realism of the synthesized avatars. While GuessTheUnseen can infer unseen-view appearances, it introduces noisy textures and multi-face ‘Janus’ artifacts, as shown in Fig. 5.
@@
full model w/o Opt. 19.78 0.9463 0.0626
@@
TABLE III: Quantitative results for ablation study. Opt. includes view selection and pose feature injection. Prog. is short for progressive training strategy.
5.3.2 In-the-wild Dataset
The in-the-wild dataset comprises various monocular dancing videos sourced from the internet. For quantitative evaluation, our method demonstrates performance comparable to the baselines for the seen parts of reconstructed humans, as shown in Tab. II. However, due to the lack of novel view references, our primary focus is on qualitative evaluation results in comparison to other methods for the unseen parts of humans. As shown in Fig. 4&5, we observe that HumanNeRF faces similar challenges to Instant-NVR. The generative networks struggle to effectively fuse sampling points for novel view rendering due to a lack of supervision. This shortcoming results in floating points and “foggy” artifacts in the rendered outputs. Additionally, the Gaussian-based methods continue to produce unrealistic “tattoo-like” appearances on the backs of synthesized avatars, highlighting its limitations in preserving overall appearance fidelity. For GuessTheUnseen, the ‘Janus’ artifacts become more pronounced, and it even fails to correctly infer the appearance of some subjects. In contrast, our method consistently demonstrates superior performance in addressing the challenges of invisible parts synthesis, excelling in both geometry and appearance reconstruction. The invisible parts of the synthesized avatars show not only enhanced geometric precision but also significantly improved appearance fidelity, with fewer artifacts and smoother textures.
@@ 5.4.1 Dual-space Optimization
Next, we evaluate the effectiveness of Dual-Space Optimization, with ablation results presented in Fig. 6. Observation optimization is crucial for reconstructing the invisible parts of the avatar, but it often encounters challenges and may not converge effectively, resulting in rough and less satisfactory appearances. In such cases, the canonical optimization step becomes essential. Leveraging the canonical space, the optimization converges more effectively, yielding smoother and more visually pleasing results. Nevertheless, observation optimization can mitigate the quality degradation issues that arise when relying solely on SDS optimization in the canonical space, as shown in Fig. 2(a),2(b)&2(c). This iterative approach highlights the importance of both observation and canonical optimization for achieving optimal reconstruction outcomes.
@@ 5.4.2 View Selection and Pose Feature Injection
We investigate the influence of View Selection and Pose Feature Injection in the following. As shown in Fig. 7, view selection filters out visible views, preserving the alignment between visible and canonical appearances, thereby reducing potential disruptions from observation optimization. Additionally, pose feature injection plays a crucial role in further enhancing dynamic appearance, allowing for the capture of finer details, especially in facial regions and cloth textures. These improvements significantly contribute to the overall fidelity and realism of the synthesized avatars.In Tab. III, we present a quantitative evaluation of all components in the ZJU-Mocap(revised) dataset, MVHumanNet dataset, and Monocap dataset. Our full model, coupled with progressive training, achieves the best results in this evaluation.
Figure 6: Ablation study about Dual-pace Optimization (a) conditioning image for SDS. (b) w/o dual space optimization. (c) w/ observation optimization only. (d) full model novel view.
Figure 7: Ablation study about View Selection and Pose Feature Injection (a) w/o both. (b) w/ view selection and w/o pose feature injection. (c) full model novel view.
5.5 Novel Poses Animation
Figure 8: Avatar animation with novel poses.
5.6 Efficiency
As shown in Fig. 1, our method, with its two-stage training process, requires approximately 3 training hours, significantly outperforming HumanNeRF, which demands about 10 hours of training on the same device. Furthermore, our method can achieve almost real-time rendering speed at 18 fps. But Instant-NVR and HumanNeRF can only render with 2 fps and 7 fps respectively.
@@ 6.1 Limitation
Since our method depends on human body fitting and foreground segmentation, artifacts may occur due to inaccuracies in these videos within the processes. Despite incorporating pose optimization to correct poses, the reconstruction of hand parts and body shape may still exhibit artifacts in certain cases, as shown in the videos. While our approach generally yields more realistic results, similar to many existing methods [15, 10, 16], it still faces challenges in accurately modeling loose attire, such as dresses, underscoring areas for potential improvement in future iterations.
@@ Appendix A Training Details
This section provides more details about the implementation and training of our method.
@@ A.1.1 Canonical Initialization
We unwrap the T-pose body onto a UV map, where each pixel stores a 3D position vector. The positional UV map, with a resolution of $(512\times 512\times 3)$, is used to initialize Gaussians in the canonical space, ensuring proper alignment with the body’s structure. Additionally, a downsampled $(128\times 128\times 3)$ version of the positional UV map serves as input to the Gaussian decoder, aiding in reconstructing and refining the 3D representation. Furthermore, we use blend
A.1.2 Training
@@
A.1.3 Pose Optimization
Our method leverages pose optimization from GaussianAvatar [16] for the In-the-wild dataset as a correction for fitted SMPL [49] pose parameters. We have omitted this functionality for the ZJU-Mocap dataset, as their ground truth pose is accurate. However, GausianAvatar keeps optimizing pose parameters for ZJU-Mocap dataset, which leads to inaccurate poses, especially for invisible parts. Please check the accompanying video results for more details.
@@ A.2.1 Dual-space Optimization
In this stage, we apply Dual-space optimization on top of visible appearance reconstruction to predict the invisible appearance. During training, each epoch is divided into three parts: 50% for given view training and 50% for Dual-space optimization. In Dual-space optimization, the weight of canonical optimization is treated as a hyperparameter, defaulting to 50%. the The fine-tuning losses are added upon $\mathcal{L}_{StageI}$. We set $\lambda_{p}=0.5$ and $\lambda_{SDS}=0.3$ initially.
A.2.2 Progressive Training
We design a progressive training strategy in this stage, gradually diminishing the weight of SDS loss. This strategy is employed to enhance further the effectiveness and efficiency of the visible appearance reconstruction. Based on this strategy, the $\lambda_{SDS}$ is reduced gradually by following:
@@ A.3 Resolution
Figure 9: Qualitative comparison results with SIFU [99] and SITH [18].
@@ Appendix B More Experiments
B.1 Comparison with Image-based Methods
@@
ZJU-Mocap(revised) ELICIT 19.23 0.9456 0.0.0689
@@
TABLE IV: Quantitative evaluation on MVHumanNet, ZJU-Mocap(revised), and Monocap datasets.
In Tab. IV, we present the quantitative evaluation results. SIFU and SITH were tested on the Monocap dataset, while ELICIT was evaluated on the ZJU-Mocap(revised) dataset. The results demonstrate that our method consistently achieves superior performance compared to the state-of-the-art approaches, underscoring its efficacy and robustness.

### v2

@@ 5.3.1 ZJU-Mocap(revised), MVHumanNet, and Monocap Datasets
These three datasets serve as the primary testbeds for our experiments due to the availability of ground truths for invisible parts. The quantitative results are presented in Tab. V, where our method surpasses all evaluation techniques in all three metrics, indicating its efficacy in reconstructing both geometry and appearance for invisible parts. Qualitatively, as shown in Fig. 4, the limitations of all compared methods become more evident when visualizing invisible parts. Methods such as GaussianAvatar, ExAvatar, and SplattingAvatar, which are based on Gaussian Splatting, exhibit noticeable artifacts and inconsistencies, including noisy textures and blank spots. Instant-NVR and HumanNeRF, due to their NeRF-based ray-shooting geometric reconstruction technique, not only struggle with appearance consistency but also suffer from geometric issues like floating artifacts and penetrating holes, diminishing the realism of the synthesized avatars. While GuessTheUnseen can infer unseen-view appearances, it introduces noisy textures and multi-face ‘Janus’ artifacts, as shown in Fig. 5.
5.3.2 In-the-wild Dataset
The in-the-wild dataset comprises various monocular dancing videos sourced from the internet. Due to the lack of novel view references, our primary focus is on qualitative evaluation results in comparison to other methods for the unseen parts of humans. As shown in Fig. 4&5, we observe that HumanNeRF faces similar challenges to Instant-NVR. The generative networks struggle to effectively fuse sampling points for novel view rendering due to a lack of supervision. This shortcoming results in floating points and “foggy” artifacts in the rendered outputs. Additionally, the Gaussian-based methods continue to produce unrealistic “tattoo-like” appearances on the backs of synthesized avatars, highlighting their limitations in preserving overall appearance fidelity. For GuessTheUnseen, the ‘Janus’ artifacts become more pronounced, and it even fails to correctly infer the appearance of some subjects. In contrast, our method consistently demonstrates superior performance in addressing the challenges of invisible parts synthesis, excelling in both geometry and appearance reconstruction. The invisible parts of the synthesized avatars show not only enhanced geometric precision but also significantly improved appearance fidelity, with fewer artifacts and smoother textures.
Figure 6: First row: Training ratio between canonical optimization and observation optimization(Canonical opt. : Observation opt.). Second row: Training ratio between Dual-space Optimization and given-view training(Dual-space : Given view).
@@ 5.3.2 In-the-wild Dataset
full model w/o View-selection. 20.03 0.9489 0.0602
full model w/o Pose-injection. 19.78 0.9473 0.0626
@@
TABLE II: Quantitative results for ablation study. Prog. is short for progressive training strategy.
@@ 5.4.1 Dual-space Optimization
Next, we evaluate the effectiveness of Dual-Space Optimization, with ablation results presented in Fig. 9. Observation optimization is crucial for reconstructing the invisible parts of the avatar, but it often encounters challenges and may not converge effectively, resulting in rough and less satisfactory appearances. In such cases, the canonical optimization step becomes essential. Leveraging the canonical space, the optimization converges more effectively, yielding smoother and more visually pleasing results. Nevertheless, observation optimization can mitigate the quality degradation issues that arise when relying solely on SDS optimization in the canonical space, as shown in Fig. 2(a),2(b)&2(c). This iterative approach highlights the importance of both observation and canonical optimization for achieving optimal results.
The balance between these two optimizations is crucial. As shown in Fig. 6 and Tab. III, the optimal ratio between canonical-space optimization and observation-space optimization is 50:50 within each training epoch. A higher proportion of canonical optimization leads to a double-face artifact, while a higher proportion of observation-space optimization results in degraded quality in the back-view appearance.
Furthermore, we continue given-view training in Stage II to preserve sharp front-view quality. As shown in Fig. 6 and Tab. III, a higher ratio of Dual-space Optimization leads to lower reconstruction quality and fidelity in the front view of the avatar. However, when the Dual-space Optimization ratio drops below 50%, the quality of the predicted back-view appearance degrades. Therefore, we empirically adopt a 50:50 training ratio to balance the quality of both views.
Dual-space. : Given view Canonical : Observation
[…]
100 : 0 19.89 0.9459 0.0626 19.53 0.9456 0.0715
75 : 25 20.56 0.9496 0.0593 19.93 0.9463 0.0698
50 : 50 21.06 0.9536 0.0551 21.06 0.9536 0.0551
25 : 75 20.67 0.9503 0.0586 20.69 0.9490 0.0596
0 : 100 19.23 0.9436 0.0667 20.06 0.9471 0.0629
TABLE III: Quantitative evaluation on training percentage ablation: Dual-space Optimization: Given view training and Canonical optimization : Observation optimization.
[…]
0% 19.80 0.9463 0.0634
25% 20.89 0.9516 0.0549
50% 21.06 0.9536 0.0551
75% 20.76 0.9503 0.0593
100% 20.16 0.9493 0.0601
TABLE IV: Quantitative evaluation on visibility ratios.
Figure 7: Qualitative comparison of results under varying visibility ratios.
((a))
((b))
((c))
((d))
Figure 8: Ablation study about View Selection (View.) and Pose Feature Injection (Pose.) (a) w/o both. (b) w/ View. and w/o Pose. (b) w/o View. and w/ Pose. (d) full model.
@@ 5.4.2 View Selection and Pose Feature Injection
We investigate the influence of View Selection and Pose Feature Injection in the following. As shown in Fig. 8, View selection filters out visible views, preserving the alignment between visible and canonical appearances, thereby reducing potential disruptions from observation optimization. Additionally, pose feature injection plays a crucial role in further enhancing dynamic appearance, allowing for the capture of finer details, especially in facial regions and cloth textures. These improvements significantly contribute to the overall fidelity and realism of the synthesized avatars. In Tab. II, we present a quantitative evaluation of all components in the ZJU-Mocap(revised) dataset, MVHumanNet dataset, and Monocap dataset. Our full model with progressive training achieves the best results in this evaluation.
In practice, we set the visibility ratio to 50% in Eq. 13 to determine whether a surface region is sufficiently observed during Stage I. This threshold is critical for balancing the learning focus between well-observed and less-visible regions in Stage II. As shown in Fig. 7, when the ratio is too low, Dual-space Optimization overfits front-view regions, producing low-fidelity artifacts and limited generalization. Conversely, setting it too high shifts the focus to poorly observed regions (e.g., side views), leading to incomplete supervision and degraded quality. Tab. IV reports that a 50% ratio provides a good trade-off, preserving high-quality reconstruction for both front and side views.
@@ 6.1 Limitation
Since our method depends on human body fitting and foreground segmentation, artifacts may occur due to inaccuracies in these videos within processes. Despite incorporating pose optimization to correct poses, the reconstruction of hand parts and body shape may still exhibit artifacts in certain cases, as shown on the left side of Fig. 10. While our approach generally yields more realistic results, similar to many existing methods [15, 10, 16], it still faces challenges in accurately modeling loose attire, such as dresses. This limitation, illustrated on the right side of Fig. 10, underscores the need for future enhancements to better capture complex clothing geometry and dynamics.
((a))
((b))
((c))
((d))
Figure 9: Ablation study about Dual-space Optimization (a) w/o Dual space Optimization. (b) w/ observation optimization only. (c) full model novel view.
Figure 10: Failure cases on hand and garments.
@@ 6.2 Conclusion
Acknowledgment
The authors would like to thank the anonymous reviewers for their constructive comments and suggestions. X. Dong is supported in part by Guangdong and Hong Kong Universities ”1+1+1” Joint Research Collaboration Scheme.The opinions expressed are solely those of the authors, and do not necessarily represent those of the funding agencies.
@@ Appendix A Training Details
This section provides more details about the implementation and training of our method. The entire pipeline was evaluated on a system equipped with two RTX 3090 GPUs using a 1920×1080 resolution video. The preprocessing stage, which includes SMPL [49] human body fitting and Sapiens [92] normal prediction, takes approximately 40 minutes in total (20 minutes each). This is followed by Stage I training, which requires around 40 minutes, and a subsequent post-processing step to generate normal maps for unseen views, taking an additional 20 minutes. Finally, Stage II training takes approximately 3 hours, bringing the total processing time to just over 4.5 hours.
@@ A.1.1 Canonical Initialization
We unwrap the T-pose body onto a UV map, where each pixel stores a 3D position vector. The positional UV map, with a resolution of $(512\times 512\times 3)$, is used to initialize Gaussians in the canonical space, ensuring proper alignment with the body’s structure. Additionally, a downsampled $(128\times 128\times 3)$ version of the positional UV map serves as input to the Gaussian decoder, aiding in reconstructing and refining the 3D representation.
A.1.2 Model Architecture
The Gaussian parameter decoder consists of an 8-layer Multi-Layer Perceptron (MLP) with intermediate layer dimensions of $(128\times 128,128\times 128,128\times 128,256\times 256,128\times 128,128\times 128,128\times 128,64\times 64)$, and includes a skip connection from the input to the fourth layer. It is followed by four separate prediction heads, each implemented as a 2-layer MLP, responsible for predicting the Gaussian position offsets $\Delta x$, normal offsets $\Delta n$, colors $c$, and scales $s$, respectively.
A.1.3 Training
@@
A.1.4 Pose Optimization
Our method leverages pose optimization from GaussianAvatar [16] for the In-the-wild dataset as a correction for fitted SMPL [49] pose parameters. We have omitted this functionality for the ZJU-Mocap dataset, as their ground truth pose is accurate. However, GaussianAvatar keeps optimizing pose parameters for the ZJU-Mocap dataset, which leads to inaccurate poses, especially for invisible parts. Please check the accompanying video results for more details.
@@ A.2.1 Dual-space Optimization
In this stage, we apply Dual-space Optimization on top of visible appearance reconstruction to predict the invisible appearance. During training, each epoch is divided into two parts: 50% for given view training and 50% for Dual-space Optimization. In Dual-space Optimization, the weight of canonical optimization is treated as a hyperparameter, defaulting to 50%. In practice, ratios between optimizations around 50:50 (e.g., 40:60 or 60:40) also yield acceptable results depending on the subject. In this work, we empirically adopt a 50:50 split as our default setting. The fine-tuning losses are added on top of $\mathcal{L}_{StageI}$. We set $\lambda_{p}=0.5$ and $\lambda_{SDS}=0.3$ initially.
A.2.2 View Selection
For view selection, we initially sample camera poses uniformly distributed around the avatar in the horizontal plane. Specifically, we fix the camera parameters with an elevation of 0 degrees and a distance of 5 units from the avatar. The azimuth angles are evenly sampled at 100 intervals to ensure full 360-degree coverage. Additionally, to enhance completeness and capture top-down information, we include one overhead view with a high elevation angle. In practice, the visibility of different parts of the human body can vary significantly due to factors such as pose, occlusion, and camera placement. As a result, the ratio between visible regions (e.g., front vs. back views) is not consistent across samples or subjects. To account for this variability, we treat the visibility ratio as a hyperparameter, with a default value of 50%, which can be tuned empirically based on the characteristics of the dataset. This allows our method to adapt to datasets with different view distributions and ensures balanced optimization of both visible and occluded regions.
A.2.3 Progressive Training
We design a progressive training strategy in this stage, gradually diminishing the weight of SDS loss. This strategy is employed to enhance further the effectiveness and efficiency of the visible appearance reconstruction. Based on this strategy, the $\lambda_{SDS}$ is reduced gradually as follows:
@@ Appendix B More Experiments
B.1 Comparison with Video-based Methods
As discussed in Section 4.2, directly applying Dual-space Optimization results in degraded reconstruction quality for seen views. To overcome this limitation, we introduce View Selection and Pose Injection. Tab. V reports the reconstruction quality for seen views, showing that our method consistently achieves performance on par with, and often surpasses, other state-of-the-art approaches. Moreover, in scenarios with significant occlusion, our method demonstrates clear advantages and see Section B.3 for more discussion.
ZJU-Mocap(revised) Monocap MVHumanNets In-the-wild
[…]
HumanNeRF 30.31 0.9642 0.0328 30.41 0.9743 0.0321 30.23 0.9756 0.0314
Instant-NVR 30.50 0.9716 0.0384 30.73 0.9763 0.0273
SplattingAvatar 28.93 0.9682 0.0326 29.36 0.9703 0.0296 28.89 0.9653 0.0306 28.28 0.9693 0.0286
ExAvatar 30.58 0.9815 0.0269 29.68 0.9732 0.0306 29.45 0.9682 0.0291 29.46 0.9709 0.0253
GaussianAvatar 29.94 0.9795 0.0210 30.98 0.9701 0.0193 28.58 0.9616 0.0237 29.96 0.9716 0.0220
GuessTheUnseen 29.36 0.9535 0.0365 28.86 0.9547 0.0326 29.35 0.9685 0.0256
Ours 29.82 0.9786 0.0219 30.56 0.9698 0.0216 28.49 0.9603 0.0259 29.76 0.9712 0.0222
TABLE V: Quantitative evaluation on seen views for ZJU-Mocap(revised), MVHumanNet, Monocap datasets, and In-the-wild dataset.
B.2 Comparison with Image-based Methods
@@
In Tab. VI, we present the quantitative evaluation results. SIFU and SITH were tested on the Monocap dataset, while ELICIT was evaluated on the ZJU-Mocap(revised) dataset. The results demonstrate that our method consistently achieves superior performance compared to the state-of-the-art approaches, underscoring its efficacy and robustness.
Figure 11: Qualitative comparison results with SIFU [101] and SITH [18].
@@
ZJU-Mocap(revised) ELICIT 19.23 0.9456 0.0689
@@
TABLE VI: Quantitative evaluation on MVHumanNet, ZJU-Mocap(revised), and Monocap datasets.
B.3 Occlusion-aware Reconstruction
Reconstructing occluded regions remains a significant challenge in full-body avatar modeling, particularly when parts of the subject are not visible in the input views. Our method addresses this issue by leveraging canonical-space priors and Dual-space Optimization (SDS) to infer and recover plausible appearance and geometry in occluded areas.
We identify three common scenarios of occlusion in video data. The first scenario involves transient occlusions, where the occluded region is visible in other frames of the video. This is the simplest case, as our method utilizes the canonical space to fuse information from well-captured frames, enabling high-fidelity reconstruction of the occluded frame without quality degradation.
The second scenario involves partial occlusion, where a small region of the body remains blocked throughout the entire video. In such cases, our method uses a visibility map to identify the occluded regions and leverages the visible body parts as guidance. It then applies Dual-space optimization to infer the texture and geometry of the occluded areas. By combining those techniques, the model is able to reconstruct visually plausible and coherent appearances for the occluded regions, as shown in Fig. 13.
The third scenario involves severe occlusion, where a large portion of the body—such as an entire leg or half of the torso—is consistently blocked throughout the video. In this case, accurate body pose fitting becomes highly challenging, as existing methods can only align to visible regions. Although one possible solution is to use Zero123 to hallucinate unseen body parts from a front-view reference image, it requires that the missing region be present in the reference. If the occluded parts are absent in both the input and reference views, our method cannot recover them. Therefore, our current pipeline is not capable of handling such extreme occlusions.
Figure 13: Demonstration of occlusion-aware reconstruction.
B.4 Efficiency
Our method, with its two-stage training process, requires approximately 3 training hours, significantly outperforming HumanNeRF, which demands about 10 hours of training on the same device. Furthermore, our method can achieve almost real-time rendering speed at 18 FPS. But Instant-NVR and HumanNeRF can only render with 2 FPS and 7 FPS respectively. In comparison to 3DGS-based approaches, our method is slower than GuessTheUnseen [26] (150 FPS) but operates at a similar speed to GaussianAvatar [16].
The training process takes 300 epochs, corresponding to approximately 60,000 iterations, which is slower compared to ExAvatar [25](10,000 iterations), SplattingAvatar [24](15,000 iterations), and GaussianAvatar [16](30,000 iterations). This increased training time is primarily due to the slower convergence of the SDS loss and the additional supervision from the normal map loss.
B.5 Novel Poses Animation
Figure 14: Avatar animations with novel poses.
Figure 15: More qualitative comparison with GuessTheUnseen [26].
Figure 16: More qualitative comparison with HumanNeRF [10], Instant-NVR [15], SplattingAvatar [24], ExAvatar [25] and GaussianAvatar [16].